OpenStudy (anonymous):

logbase3(2x+4)-logbase3(x)=4

OpenStudy (anonymous):

are you stii there

OpenStudy (anonymous):

\[\log_{3}(2x+4) - log_{3}(x) = 4\] \[\log_{3}(2x+4)/(x) = 4\] \[\log_{3}(2x+4)/(x) = log_{3} 81\] \[(2x+4)/(x) = 81\] \[(2x+4) = 81x\] \[79x = 4\] \[x = 4/71\] I hope I did not make a mistake

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

x=79/4

OpenStudy (anonymous):

great work kosh as usual!

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

x=4/79

OpenStudy (anonymous):

great kushhhh u always did better